OpenStudy (anonymous):

what is the area of triangle whose two sides are 5 and one of the sides is 4 ??

OpenStudy (anonymous):

u can use herons formula for that...

Parth (parthkohli):

Heron was pretty cool for creating a formula like that :)

OpenStudy (anonymous):

i don't know that...i know this is an isosceles triangle :(

OpenStudy (anonymous):

s=(5+5+4)/2 = 7 so Area = \[\sqrt{s(s-a)(s-b)(s-c)}\]

OpenStudy (anonymous):

thanks...but this formula is so long and boring ...sorry :)

hartnn (hartnn):

lol!

hartnn (hartnn):

no other option i guess..

OpenStudy (anonymous):

kkk u know dat area of triangle = 1/2 * base * height

OpenStudy (anonymous):

1/2 b*h where height you get from pythogoras theorem and base is 4

hartnn (hartnn):

\(\sqrt{s(s-a)(s-b)(s-c)}\) was better, seriously.

OpenStudy (anonymous):

so by constructing a perpendicular on the base of 4 cm and that perpendicular will bisect the side

OpenStudy (anonymous):

i am confused lol

OpenStudy (ghazi):

area = 2* (0.5*2*h) and h= sqrt 16-4

OpenStudy (ghazi):

\[h= 2\sqrt2\]

OpenStudy (anonymous):

UnkleRhaukus: Your response is incorrect. Please reread the question.

OpenStudy (anonymous):